# Inference-Time Selective Debiasing

### v1

@@ 4.2 Results
Table 1 compares fairness, accuracy, DTO, and FF-score for various at-training and pre-processing debiasing methods, post-processing debiasing methods, and for selective debiasing based on LEACE with SR and the proposed KL-based bias quantification score. Here, we show results only for the optimal selection percentage. The full results with various selection percentages are presented in Appendix B. The optimal percentage was chosen over the first 15% of the validation set. “LEACE-last” represents LEACE applied to the outputs of the last hidden layer of the classifier, while “LEACE-cls” is LEACE applied to each linear layer of the classification head of the model. For comparison, we also provide the results with selective debiasing using INLP. The post-processing methods were trained on only 20% of the training set.
In the majority of cases, the best results are unsurprisingly achieved by at-training and pre-processing debiasing techniques, as these methods retrain the models from scratch on the full training data. Nevertheless, the proposed selective debiasing approach based on LEACE substantially enhances the results of inference-time debiasing using post-processing techniques in terms of metrics that take into account both fairness and performance: FF-score and DTO. It is also competitive with at-training and pre-processing techniques. For LEACE-cls with KL selection, selective debiasing even outperforms these methods on MOJI-balanced. The results in Tables 10, 11 and 12 also show that selective debiasing consistently outperforms standard inference-time debiasing in terms of FF-score across various percentages.
LEACE-cls generally achieves better fairness than LEACE-last and slightly better joint fairness–performance, in terms of DTO and FF-score. INLP-based selective debiasing improves the FF-score only on MOJI-balanced, while on other datasets, it is consistent with the base inference-time debiasing method. INLP-based approaches overall fall behind the corresponding LEACE-based techniques.
When comparing the results of the proposed bias quantification method based on the KL distance with SR, we can see that our method notably outperforms SR on MOJI datasets and is on par with SR on BIOS-2. We further explore other distance-based bias quantification methods (Euclidean and cosine distances), which are presented in Appendix C. Results in Tables 10, 11 and 12 show that in most cases, selection by KL works comparably or better than other distance-based measures. Moreover, KL scores are easier to compute than distance-based scores.
@@ 5 Conclusion and Future Work
We proposed selective debiasing – a new simple inference-time safety mechanism for increasing model performance and fairness. We showed that it is helpful in the case that re-training a model from scratch for better fairness is prohibitive or there is no access to full training data. Additionally, for the selection of problematic predictions, we suggest a bias quantification approach based on KL distance that achieves better results than the standard UQ method. The proposed mechanism fills the gap for efficient techniques that can be applied at inference time and opens the door for safer ML-based systems. In future work, we aim to investigate a deeper integration between UQ and debiasing methods.
@@
Appendix A Training Setup and Hyperparameters
To find an optimal set of hyperparameters, we conducted a grid search on the validation set. We used accuracy as an optimization target for standard models, and DTO for models with debiasing. The grid and optimal parameters for the standard models are described in Table 2. For each debiasing method, we tuned the method’s parameters and kept training parameters of the base model – the grid and optimal values for debiasing methods presented in Table 3. The training was conducted on a cluster with Nvidia V100 GPUs. An approximate number of GPU-hours spent during the experiments is presented in Table Table 4.
@@ Appendix A Training Setup and Hyperparameters
Appendix B Selective Debiasing on Various Percentages
To check how stable the proposed methods are, we compare selective debiasing results over 5%, 10%, and 15% of selection for random, SR, and KL scores. The results are presented in Tables 6, 7 and 8. The optimal percentage selected on the validation set from values from 1% to 15%; results for each dataset-method pair in Table 9. In general, optimal scores are better or comparable with results on various percentages, which allows us to use this approach to detect the optimal percentage of selection. Table 5 shows the performance of selective debiasing and post-processing debiasing methods trained on a full training set. As one can see, the performance on the full set is comparable with the results on only 20% from Table 1.
@@ Appendix B Selective Debiasing on Various Percentages
Table 5: Comparison of debiasing methods and selective debiasing; the post-processing methods trained on full training set. The best results in the group are in bold, and the best results overall are underlined.
@@
Appendix C Comparison with Other Distances
We also conducted additional experiments to compare how proposed selection strategies differ from other similarity measures. Here we consider several measures, calculated over output from the last hidden layer of the model, and compare it with SR and KL strategies. The results are presented in Tables 10, 11 and 12. In most cases, selection by KL works comparably or better than the best-performing distance-based measure. Moreover, KL scores are easier to compute than distance-based scores. However, in some cases, cosine distance could serve as a replacement for the KL score due to its similar performance.
@@ Appendix C Comparison with Other Distances
Appendix D Datasets Statistics
The synthetic dataset was generated as a random 2 classes classification task using make_classification function from Scikit-learn library Pedregosa et al. (2011) with following parameters: n_features=10, n_informative=5, n_clusters_per_class=2, random_state=42, n_redundant=2. The protected attribute for the synthetic dataset is designed as a condition over the first informative feature and equals 1 if this feature is greater than 0, and 0 otherwise. The overall statistics for each dataset are presented in Table 13. Tables 14 and 15 shows the joint distribution of target variable and protected attributes.
Dataset
Num. of
classes/attributes
Protected
attribute
Train/Val/Test
Synthetic 2/2 Geometric 6k/2k/2k
Moji (balanced) 2/2 Race 100k/8k/8k
Moji (imbalanced) 2/2 Race 100k/5k/5k
Bios-2 (imbalanced) 2/2 Gender 21k/3k/8k
Bios-2 (balanced) 2/2 Gender 21k/1k/2k
Table 13: Dataset statistics.
Split Gender Profession
Nurse Surgeon Total
Train Female 53.34 5.74 59.08
Male 5.50 35.42 40.92
All 58.84 41.16 100.00
Val Female 53.32 5.08 58.40
Male 5.52 36.08 41.60
All 58.83 41.17 100.00
Test Female 53.82 7.51 61.33
Male 5.01 33.66 38.67
All 58.83 41.17 100.00
Val (balanced) Female 26.02 23.98 50.00
Male 26.02 23.98 50.00
All 52.05 47.95 100.00
Test (balanced) Female 20.02 29.98 50.00
Male 20.02 29.98 50.00
All 40.04 59.96 100.00
Table 14: Joint distribution for BIOS-2 dataset.
Split Ethnicity Target
Sad Happy Total
Train SA 40.00 10.00 50.00
AA 10.00 40.00 50.00
All 50.00 50.00 100.00
Val SA 40.02 9.98 50.00
AA 9.98 40.02 50.00
All 50.00 50.00 100.00
Test SA 40.02 9.99 50.01
AA 9.99 40.00 49.99
All 50.01 49.99 100.00
Val (balanced) SA 25.00 25.00 50.00
AA 25.00 25.00 50.00
All 50.00 50.00 100.00
Test (balanced) SA 25.01 25.01 50.01
AA 24.99 24.99 49.99
All 50.00 50.00 100.00
Table 15: Joint distribution for MOJI dataset.
Appendix E Equal Opportunity
There are numerous amount of group fairness definitions; to avoid any mismatches, we are presenting the step-by-step process of equal opportunity criterion calculation. This criterion is based on recall values, or true positive rate (TPR) for each class and protected group.
  • TPR (recall) for each protected group defined as follows:
    $TPR=\frac{TP}{TP+FN},$ (5)
    where $TP$, $FN$ is true positives and false negatives for specific group.
  • After we calculate TPR-gap:
    $\delta=\sqrt{\frac{1}{C}\sum_{c}(\sum_{g}|TPR_{c,g}-\overline{TPR_{c}}|)^{2}},$ (6)
  • Finally, we calculate fairness with the following equation:
    $Fairness=100\cdot(1-\delta).$ (7)
Appendix F Fairness and UQ Oracles
In this section, we describe in detail oracle strategies for fairness and accuracy. For both strategies, we assume access to the ground-truth labels, while for fairness oracle we also use protected attributes. Accuracy oracle built as follows – we find all erroneously classified instances and replace predictions on these instances with ground-truth labels while keeping all other predictions unchanged. This oracle shows the best possible UQ strategy that allows the detection of all erroneous predictions and gives the maximal increase in accuracy. The same idea is behind fairness oracle, but instead of accuracy, we used fairness as a target metric. For fairness, we first replace predictions for instances, which gives the maximal increase in fairness. These predictions were chosen greedily from the erroneous ones. To measure the quality of these oracle strategies and to compare it with other scores, we calculated several metrics: FR-AUC, Acc-AUC, and FF-score-AUC. Each corresponds to the area under the target metric-rejection curve, where the target metric is fairness, accuracy, or FF-score; the area under the curve is calculated on binarized over 100 points target metric values.

### v3

@@ 4.2 Results
Table 1 presents results for various at-training and pre-processing debiasing methods, post-processing debiasing methods, selective debiasing based on LEACE with SR, and selective debiasing using the proposed KL-based bias quantification score. Here, we show results only for the threshold that gives an optimal selection percentage. The full results with various selection percentages are presented in Appendix F. The results for selective debiasing using INLP are provided in Appendix F.
In the majority of cases, the best results are unsurprisingly achieved by at-training and pre-processing debiasing techniques, as these methods retrain the models from scratch on the full training data. Nevertheless, the proposed selective debiasing approach based on LEACE substantially enhances the results of inference-time debiasing using post-processing techniques in terms of metrics that take into account both fairness and performance: FF-score and DTO. Inference time debiasing becomes competitive with at-training and pre-processing techniques. For LEACE-cls with KL selection, selective debiasing even outperforms these methods on MOJI-balanced. The results in Tables 15, 16 and 17 also show that selective debiasing consistently outperforms standard inference-time debiasing in terms of FF-score.
LEACE-cls generally achieves better fairness than LEACE-last and slightly better joint fairness–performance in terms of DTO and FF-score.
When comparing the results of the proposed bias quantification method based on the KL distance with SR, we can see that our method notably outperforms SR on the MOJI datasets and is on par with SR on BIOS-2. We further explore other distance-based bias quantification methods (Euclidean and cosine distances) in Appendix G. Results in Tables 15, 16 and 17 show that in most cases, selection by KL works comparably or better than other distance-based measures. Moreover, KL scores are easier to compute than distance-based scores.
@@ 5 Conclusion and Future Work
We proposed selective debiasing – a new simple inference-time safety mechanism for increasing model performance and fairness. We showed that it is helpful in the case when re-training a model from scratch for better fairness is prohibitive or there is no access to full training data. Additionally, for the selection of problematic predictions, we suggest a bias quantification approach based on KL divergence that achieves better results than the standard UQ method. The proposed mechanism fills the gap for efficient techniques that can be applied at inference time and opens the door for safer ML-based systems. In future work, we aim to investigate a deeper integration between UQ and debiasing methods.
@@
Acknowledgments
We appreciate the anonymous reviewers for their valuable suggestions that helped enhance this paper. This research was supported in part through the computational resources of HPC facilities at HSE University.
Appendix A LEAst-Squares Concept Erasure
LEACE removes information about a concept $Z$ from the representation space $X$. To formally describe LEACE, we firstly introduce the following notions. Let $\mathbf{x}\in X$ be an instance from $X$ (e.g. embedding from the last layer in the case of LEACE-last), $\mathbf{\Sigma_{XX}}$ is the covariance matrix for $X$, $\mathbf{\Sigma_{XZ}}$ is the covariance matrix between $X$ and $Z$, and $W_{\perp}$ stands for the pseudoinverse of the matrix $W$. The $\mathbf{W}$ and $\mathbf{P}_{\mathbf{W}\mathbf{\Sigma}_{XZ}}$ defined as follows:
$\mathbf{W}=(\mathbf{\Sigma_{XX}}^{1/2})_{\perp},$ (5)
$\mathbf{P}_{\mathbf{W}\mathbf{\Sigma}_{XZ}}=(\mathbf{W}\mathbf{\Sigma_{XZ}})(\mathbf{W}\mathbf{\Sigma_{XZ}})_{\perp}.$ (6)
Then the final LEACE transformation is defined as follows:
$\hat{\mathbf{y}}(\mathbf{x})=\mathbf{x}-\mathbf{W}_{\perp}\cdot\mathbf{P}_{\mathbf{W}\mathbf{\Sigma}_{XZ}}\cdot\mathbf{W}(\mathbf{x}-\mathbb{E}[X])$ (7)
Appendix B Fairness and UQ Oracles
In this section, we describe in detail oracle strategies for fairness and accuracy. For both strategies, we assume access to the ground-truth labels, while for fairness oracle we also use protected attributes. Accuracy oracle is built as follows – we find all erroneously classified instances and replace predictions on these instances with ground-truth labels while keeping all other predictions unchanged. This oracle shows the best possible UQ strategy that allows the detection of all erroneous predictions and gives the maximal increase in accuracy. The same idea is behind fairness oracle, but instead of accuracy, we use fairness as a target metric. For fairness, we first replace predictions for instances, which gives the maximal increase in fairness. These predictions are chosen greedily from the erroneous ones. To measure the quality of these oracle strategies and to compare them with other scores, we calculated several metrics: FR-AUC, Acc-AUC, and FF-score-AUC. Each corresponds to the area under the target metric-rejection curve, where the target metric is fairness, accuracy, or FF-score; the area under the curve is calculated on binarized over 100 points target metric values.
Appendix C Datasets Statistics
The synthetic dataset was generated as a random 2 classes classification task using make_classification function from Scikit-learn library Pedregosa et al. (2011) with the following parameters: n_features=10, n_informative=5, n_clusters_per_class=2, random_state=42, n_redundant=2. The protected attribute for the synthetic dataset is designed as a condition over the first informative feature and equals 1 if this feature is greater than 0, and 0 otherwise. The overall statistics for each dataset are presented in Table 2. Tables 3 and 4 shows the joint distribution of the target variable and protected attributes.
Dataset
Num. of
classes/attributes
Protected
attribute
Train/Val/Test
Synthetic 2/2 Geometric 6k/2k/2k
Moji (balanced) 2/2 Race 100k/8k/8k
Moji (imbalanced) 2/2 Race 100k/5k/5k
Bios-2 (imbalanced) 2/2 Gender 21k/3k/8k
Bios-2 (balanced) 2/2 Gender 21k/1k/2k
Table 2: Dataset statistics.
Split Gender Profession
Nurse Surgeon Total
Train Female 53.34 5.74 59.08
Male 5.50 35.42 40.92
All 58.84 41.16 100.00
Val Female 53.32 5.08 58.40
Male 5.52 36.08 41.60
All 58.83 41.17 100.00
Test Female 53.82 7.51 61.33
Male 5.01 33.66 38.67
All 58.83 41.17 100.00
Val (balanced) Female 26.02 23.98 50.00
Male 26.02 23.98 50.00
All 52.05 47.95 100.00
Test (balanced) Female 20.02 29.98 50.00
Male 20.02 29.98 50.00
All 40.04 59.96 100.00
Table 3: Joint distribution for the BIOS-2 dataset.
Split Ethnicity Target
Sad Happy Total
Train SA 40.00 10.00 50.00
AA 10.00 40.00 50.00
All 50.00 50.00 100.00
Val SA 40.02 9.98 50.00
AA 9.98 40.02 50.00
All 50.00 50.00 100.00
Test SA 40.02 9.99 50.01
AA 9.99 40.00 49.99
All 50.01 49.99 100.00
Val (balanced) SA 25.00 25.00 50.00
AA 25.00 25.00 50.00
All 50.00 50.00 100.00
Test (balanced) SA 25.01 25.01 50.01
AA 24.99 24.99 49.99
All 50.00 50.00 100.00
Table 4: Joint distribution for the MOJI dataset.
Appendix D Training Setup and Hyperparameters
To find an optimal set of hyperparameters, we conducted a grid search on the validation set. We used accuracy as an optimization target for standard models, and DTO for models with debiasing. The grid and optimal parameters for the standard models are described in Table 5. For each debiasing method, we tuned the method’s parameters and kept the training parameters of the base model – the grid and optimal values for debiasing methods presented in Table 6. The training was conducted on a cluster with Nvidia V100 GPUs. An approximate number of GPU hours spent during the experiments is presented in Table 7.
@@ Appendix D Training Setup and Hyperparameters
Appendix E Equal Opportunity
There are a numerous amount of group fairness definitions; to avoid any mismatches, we are presenting the step-by-step process of equal opportunity criterion calculation. This criterion is based on recall values, or true positive rates (TPR) for each class and protected group.
  • TPR (recall) for each protected group defined as follows:
    $TPR=\frac{TP}{TP+FN},$ (8)
    where $TP$, $FN$ – is true positives and false negatives for specific group.
  • After we calculate TPR-gap:
    $\delta=\sqrt{\frac{1}{C}\sum_{c}\sum_{g}|TPR_{c,g}-\overline{TPR_{c}}|^{2}},$ (9)
  • Finally, we calculate fairness with the following equation:
    $Fairness=100\cdot(1-\delta).$ (10)
Appendix F Additional Experiments
To check how stable the proposed methods are, we compare selective debiasing results over 5%, 10%, and 15% of selection for random, SR, and KL scores. The results are presented in Tables 9, 10 and 11. The optimal percentage selected on the validation set from values from 1% to 15%; results for each dataset-method pair in Tables 12 and 13. In general, optimal scores are better or comparable with results on various percentages, which allows us to use this approach to detect the optimal percentage of selection.
Table 8 shows the performance of selective debiasing and post-processing debiasing methods trained on a full training set. As one can see, the performance on the full set is comparable with the results on only 20% from Table 1.
The results for selective debiasing with INLP trained on 20% of data are presented in Table 14. INLP-based selective debiasing improves the FF-score only on MOJI-balanced, while on other datasets, it is consistent with the base inference-time debiasing method. INLP-based approaches overall fall behind the corresponding LEACE-based techniques.
Debiasing method type No debiasing At-training Pre-processing Post-processing & Selective
@@ Appendix F Additional Experiments
Table 8: Comparison of debiasing methods and selective debiasing; the post-processing methods trained on full training set. The best results in the group are in bold, and the best results overall are underlined. The gray color corresponds to the results with p-value > 0.05 with respect to standard model.
@@
MOJI imbalanced 15 15 14 15 15 15 15 14 15
MOJI balanced 15 15 15 15 15 15 15 13 15
BIOS-2 imbalanced 1 6 6 1 6 4 15 6 14
BIOS-2 balanced 1 11 7 7 12 5 8 15 5
Dataset LEACE-last LEACE-cls INLP
Random SR KL Random SR KL Random SR KL
@@
Table 13: Optimal selection percentages for various debiasing methods, the post-processing methods trained on full training set.
Debiasing method type No debiasing At-training Pre-processing Post-processing & Selective
Dataset Metric Standard Adv DAdv BTEO BTJ INLP
INLP+SR,
opt. perc.
INLP+KL,
opt. perc.
MOJI imbalanced Fairness $\uparrow$ 61.8$\pm$0.7 73.7$\pm$0.6 73.4$\pm$0.4 75.2$\pm$0.6 74.8$\pm$0.6 77.3$\pm$7.3 63.5$\pm$1.1 70.5$\pm$2.4
Accuracy $\uparrow$ 79.1$\pm$0.7 72.0$\pm$0.7 72.4$\pm$0.5 73.6$\pm$0.6 73.2$\pm$0.4 68.4$\pm$6.8 78.0$\pm$0.6 73.5$\pm$1.4
DTO $\downarrow$ 43.6$\pm$0.6 38.4$\pm$0.5 38.3$\pm$0.4 36.2$\pm$0.1 36.7$\pm$0.4 40.0$\pm$3.5 42.6$\pm$0.8 39.7$\pm$1.8
FF-score $\uparrow$ 69.4$\pm$0.4 72.8$\pm$0.4 72.9$\pm$0.3 74.4$\pm$0.1 74.0$\pm$0.3 71.9$\pm$2.2 70.0$\pm$0.6 71.9$\pm$1.3
MOJI balanced Fairness $\uparrow$ 69.5$\pm$0.2 83.8$\pm$0.8 84.7$\pm$1.5 85.5$\pm$0.5 85.6$\pm$0.6 85.8$\pm$8.3 71.7$\pm$0.6 77.9$\pm$4.4
Accuracy $\uparrow$ 71.9$\pm$0.4 74.0$\pm$0.4 74.1$\pm$0.6 74.8$\pm$0.3 74.5$\pm$0.4 63.0$\pm$6.9 71.9$\pm$0.4 68.5$\pm$2.0
DTO $\downarrow$ 41.5$\pm$0.4 30.7$\pm$0.7 30.1$\pm$0.7 29.0$\pm$0.1 29.3$\pm$0.4 40.9$\pm$4.4 39.9$\pm$0.6 38.8$\pm$1.2
FF-score $\uparrow$ 70.7$\pm$0.3 78.6$\pm$0.5 79.1$\pm$0.6 79.8$\pm$0.1 79.6$\pm$0.3 71.9$\pm$3.2 71.8$\pm$0.4 72.8$\pm$1.0
BIOS-2 imbalanced Fairness $\uparrow$ 90.4$\pm$0.8 97.2$\pm$0.8 96.4$\pm$0.4 95.8$\pm$1.0 96.6$\pm$0.8 92.0$\pm$1.6 91.7$\pm$1.4 91.9$\pm$1.7
Accuracy $\uparrow$ 96.7$\pm$0.1 94.8$\pm$0.4 95.0$\pm$0.3 95.2$\pm$0.3 95.0$\pm$0.5 95.8$\pm$0.6 96.2$\pm$0.3 95.8$\pm$0.6
DTO $\downarrow$ 10.1$\pm$0.7 5.9$\pm$0.2 6.2$\pm$0.2 6.5$\pm$0.6 6.1$\pm$0.3 9.1$\pm$1.3 9.1$\pm$1.3 9.2$\pm$1.4
FF-score $\uparrow$ 93.5$\pm$0.4 96.0$\pm$0.2 95.7$\pm$0.1 95.5$\pm$0.4 95.8$\pm$0.2 93.8$\pm$0.7 93.9$\pm$0.7 93.8$\pm$0.8
BIOS-2 balanced Fairness $\uparrow$ 89.7$\pm$0.6 97.8$\pm$0.8 98.0$\pm$0.8 95.9$\pm$0.8 96.4$\pm$0.3 91.8$\pm$2.0 91.6$\pm$1.9 91.3$\pm$1.9
Accuracy $\uparrow$ 92.4$\pm$0.3 91.9$\pm$0.6 91.9$\pm$1.5 92.6$\pm$0.5 92.9$\pm$0.6 91.9$\pm$1.2 92.2$\pm$1.0 91.9$\pm$1.2
DTO $\downarrow$ 12.8$\pm$0.6 8.5$\pm$0.4 8.4$\pm$1.4 8.5$\pm$0.2 8.0$\pm$0.6 11.7$\pm$1.7 11.5$\pm$1.7 12.0$\pm$1.6
FF-score $\uparrow$ 91.1$\pm$0.4 94.7$\pm$0.1 94.9$\pm$0.7 94.2$\pm$0.2 94.6$\pm$0.3 91.8$\pm$1.1 91.9$\pm$1.2 91.6$\pm$1.1
Table 14: Comparison of debiasing methods and selective debiasing using INLP. The best results in the group are in bold, and the best results overall are underlined. The results averaged over 5 random seeds. The gray color corresponds to the results with p-value > 0.05 with respect to standard model.
Appendix G Comparison with other Distances
We also conducted additional experiments to compare how proposed selection strategies differ from other similarity measures. Here, we consider several measures, calculated over the output from the last hidden layer of the model, and compare them with SR and KL strategies. The results are presented in Tables 15, 16 and 17. In most cases, selection by KL works comparably or better than the best-performing distance-based measure. Moreover, KL scores are easier to compute than distance-based scores. However, in some cases, cosine distance could serve as a replacement for the KL score due to its similar performance.
@@ Appendix G Comparison with other Distances
Appendix H Computational Efficiency
To estimate the computational efficiency of selective debiasing, we calculated the inference time of the standard model and the model with selective debiasing. The results are presented in Tables 18 and 19. Table 18 shows the inference time of models averaged for 10 runs, while Table 19 presents computational overhead for each debiasing method. The computational overhead is calculated as follows:
$CompOverhead=100\cdot\left(\frac{T_{selective}}{T_{standard}}-1\right),$ (11)
where $T$ is the summary inference time of the debiasing method for all datasets. These experiments were conducted on one Nvidia H100 GPU. The proposed selective debiasing approach does not introduce much computational overhead – for LEACE-last and LEACE-cls it is less than 1%.
Table 20 shows a detailed comparison of debiasing methods. As one can see, at-training and pre-processing debiasing methods require a training model from scratch, while post-processing methods with selective debiasing do not require this. Hence, post-processing methods are especially beneficial when the full dataset or the model is unavailable, while selective debiasing allows for increasing the overall performance of these methods. On the other hand, there is some computational overhead for post-processing methods compared to other ones. However, this overhead is negligible in most cases.
Dataset LEACE-last LEACE-cls INLP
Selective Standard Selective Standard Selective Standard
MOJI imbalanced 3.738$\pm$0.011 3.737$\pm$0.020 3.762$\pm$0.009 3.749$\pm$0.035 3.775$\pm$0.008 3.730$\pm$0.008
MOJI balanced 5.978$\pm$0.023 5.96$\pm$0.014 6.008$\pm$0.014 5.971$\pm$0.024 6.053$\pm$0.017 5.974$\pm$0.016
BIOS-2 imbalanced 6.064$\pm$0.018 6.059$\pm$0.033 6.090$\pm$0.018 6.049$\pm$0.013 6.116$\pm$0.022 6.051$\pm$0.022
BIOS-2 balanced 1.526$\pm$0.007 1.525$\pm$0.006 1.544$\pm$0.024 1.527$\pm$0.025 1.542$\pm$0.004 1.525$\pm$0.004
Table 18: Inference time of standard model and model with applied selective debiasing (in seconds, averaged for 10 runs).
LEACE-last LEACE-cls INLP
Overhead, % 0.14 0.62 1.19
Table 19: The computational overhead of selective debiasing for various methods.
Debiasing method type Base At-training Pre-processing Selective
Debiasing method Standard Adv DAdv BTEO BTJ
LEACE-last
selective
LEACE-cls
selective
INLP
selective
Require model retraining
from Standard model
At-training method
Pre-processing method
Post-processing method
Inference speed
(relative to Standard model)
1.000 1.000 1.000 1.000 1.000 1.001 1.006 1.012
Table 20: Debiasing methods comparison. At-training and pre-processing debiasing methods can have the same inference speed, but require model training from scratch, which is impossible in some cases.